kindly do it as written form on word

1 Market Equilibrium [60 marks]

Suppose the demand and supply relations of wheat in an economy are given by:

D: Q = 1500 − 6P + aY

S: Q = −270 + 9P

where Q denotes the quantity of wheat measured in million of bushels per year, P is the price of wheat measured in thousand of £ per million bushels; Y is average income measured in thousand £ per year.

2. Let a = 0.1. Suppose that average income equals the sum of all digits in your student number (for example, for student number 123456789, Y = 45). Use Excel to draw a chart that illustrates the demand and supply relations and how these relations change if average income doubles. Use Solver to compute the exact values of equilibrium quantity and price in both cases, i.e. when income equals the sum of all digits in your student number and when income is double that amount. Comment on the changes to the supply and demand curve as a result of the doubling of income and the income effect on the equilibrium quantity and price.[25 marks] my student number is 100241246 it is required to compute value of Y

1Be careful how you label your axes!
Answer #1

Answer:

Given demand function

Q=1500-6P+aY

Supply function

Q=-270+9P

a=0.1

Y=1+0+0+2+4+1+2+4+6=20 thousand

So demand function Q=1500-6P+0.1*20=1500-6P+2

Q=1502-6P

Demand Supply Curve 250.00 Demand Curve Supply Curve L 150.00 50.00 0.00 0 200 400 600 800 1000 1200 1400 Quantity

For Equilibrium Demand = Supply

Qd=Qs

1502-6P=-270+9P

P=118.13 thousands per million bushels

Q=-270+9*118.13=793.2 million bushels

When Income doubles Y=2*20=40 thousands

So demand function Q=1500-6P+0.1*40=1504-6P

Q=1504-6P

Demand Supply Curve 250.00 Demand Curve Supply Curve L 150.00 50.00 0.00 0 200 400 600 800 1000 1200 1400 Quantity

For Equilibrium Demand = Supply

Qd=Qs

1504-6P=-270+9P

P=118.267 thousands per million bushels

Q=-270+9*118.267=794.4 million bushels

When income double then supply will remain same because supply curve is independent of income.whereas demand curve is shift slightly towards right because income has slightly impact on total demand because of income multiplier a=0.1.

Equilibrium price (from 118.13 to 118.267) and output ( from 793.2 to 794.4) will also increase slightly because income has slightly impact on total demand function because of income multiplier a=0.1.